A metallic element has a density of 7.15 g cm-3
, a lattice constant of 2.880 Å and
an atomic weight of 51.9961 u. Predict the lattice crystal structure for this element
1
Expert's answer
2020-02-03T14:25:03-0500

volume of the lattice = "a^3" = "(2.880\times10^{-8})^3 cm^3 =" "23.9*10^{-24} cm^3"

let the number of atoms per unit cell be "x"

then mass of one unit cell "= density\times volume =" "7.15\times23.9\times10^{-24}=17*10^{-23} g"

number of unit cell = "\dfrac{mass \ of \ one \ unit \ cell}{mass \ of \ one \ atom }" "=\dfrac{17*10^{-23}}{1.66*10^{-24}*52}=2"


From this we can say that the element has body centered cubic crystal